OpenStudy (anonymous):

Factor: 4x2 + 12x + 9

OpenStudy (anonymous):

check to see if there is a gcf in this case its 1

OpenStudy (anonymous):

you just look at the factors of 4 and 9 and see if any combination of them will get you 12

OpenStudy (anonymous):

4 and 3 will get me 12 but 9 is prime so what do I do now ?

OpenStudy (anonymous):

so you would use the Ac method or Decomposition a=4x^2,c=9 36 what number adds up to 12 and when multiplied =36 in this case its 6 and 6 so it would be like this (4x^2+6x) ( 6x+9) 2x(2x+3) 3(2x+3)

OpenStudy (anonymous):

my bad sorry mistake

OpenStudy (anonymous):

Actually, you can use the identity \((a+b)^2 = a^2 + 2ab+b^2\) to do it. 4x^2 + 12x + 9 = (2x)^2 + 2(2x)(3) + 3^2 = ...?

OpenStudy (anonymous):

can any1 answer my questions

OpenStudy (anonymous):

@RolyPoly its only true in this case

OpenStudy (anonymous):

Yes. But if it is applicable in this case, why not use this identity?

OpenStudy (anonymous):

im trying to finsih already this is the last test for me

OpenStudy (anonymous):

Factor: 4x^2 + 12x + 9 A (2x + 3)2 B (4x + 9)(x + 1) C (2x + 1)(2x + 9) D (2x + 3)(2x - 3)

OpenStudy (anonymous):

@SPC Which part you don't understand?

OpenStudy (anonymous):

I dont understand how you got your answer.

OpenStudy (anonymous):

A

OpenStudy (anonymous):

How to Factor step by step: x^2 - 28x + 75 A (x - 14)(x - 14) B (x - 5)(x - 15) C (x - 3)(x - 25)

OpenStudy (anonymous):

4x^2 + 12x + 9 Consider the first term, 4x^2 = \(2^2x^2\) = \((2x)^2\) Consider the second term, 12x = 2(2x)(3) Consider the last term, 9 = 3^2 Now, think about the identity for perfect square, \((a+b)^2 = a^2 + 2ab+b^2\) Match the corresponding a and b to get your answer.